To evaluate the following hypotheses

H0 :p = 0.3

HA :p ≠ 0.3


1
Expert's answer
2021-06-18T10:27:26-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.3"

"H_1:p\not=0.3"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

"=\dfrac{0.36-0.3}{\sqrt{\dfrac{0.3(1-0.3)}{50}}}\approx0.92582"

Since it is observed that "|z|=0.92582<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.3," at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=2P(Z<-0.92582)=0.35454," and since "p=0.35454>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.3," at the "\alpha=0.05" significance level.